deathscarewolf:

Elliot has been running a lawn care business since 2000. He cuts grass, trims, and weed whacks yards for his customers throughout the season. Each year, he has increased his fee by the same amount. The table shows what Elliot charged each customer for two given years of his business: (pic will be in comments give me a sec for it) A. What is the rate of change and initial value for Elliot's business? How do you know? B. Write an equation in slope-intercept form to represent the fees that Elliot charges each year.

deathscarewolf:

sophia333:

https://brainly.com/question/1556676 Heres something that might help :)

deathscarewolf:

my school blocked brainly

sophia333:

Oh lemme screenshot

sophia333:

Rate of change = (1350 - 750)/(2010 - 2000) = 600/10 = $60 Initial value = $750 Equation is y = 60x + 750; where y is the amount charged to customer and x is the number of years from 2000.

deathscarewolf:

alright

Deku:

the answer would be A correct me if you think im wrong

deathscarewolf:

deku a is a question not a awnser

Deku:

welp lol

Deku:

Initial value = value when he started ( in 2000) =$ 750 Each year he add a constants to that value let y be the amount he get after t years a amount increase each year y = 750 + at you're given that in 2010 his free = 1350 t = 2010 - 2000 = 10 year y = 1350 y = 750 + at 1350 = 750 + 10 a 600 = 10a a = 60 y = 750 + 60t ( his fee as a function of time) In slope form y = 60t + 750 so he increases his fee by 60 $ a year

deathscarewolf:

thx

Deku:

np